OpenStudy (anonymous):

plz i need help Tell whether the given ordered pairs satisfy a linear function. Explain. x -1 0 1 2 3 y -3 -2 -1 0 1

OpenStudy (aravindg):

if its a linear fn y=kx y/x =k =constant

OpenStudy (aravindg):

see whether this holds for each value of x an y

hartnn (hartnn):

if its a linear fn * y=mx+c

hartnn (hartnn):

if you notice, if you subtract 2 from x, you get y so, y = x-2